# The Dark Side of NFTs: A Large-Scale Empirical Study of Wash Trading

### v1

@@ Abstract
NFTs (Non-Fungible Tokens) have seen significant growth since they first captured public attention in 2021. However, the NFT market is plagued by fake transactions and economic bubbles, e.g., NFT wash trading. Wash trading typically refers to a transaction involving the same person or two colluding individuals, and has become a major threat to the NFT ecosystem. Previous studies only detect NFT wash trading from the financial aspect, while the real-world wash trading cases are much more complicated (e.g., not aiming at inflating the market value). There is still a lack of multi-dimension analysis to better understand NFT wash trading. Therefore, we present the most comprehensive study of NFT wash trading, analyzing 8,717,031 transfer events and 3,830,141 sale events from 2,701,883 NFTs. We first optimize the dataset collected via the OpenSea API. Next, we identify three types of NFT wash trading and propose identification algorithms. Our experimental results reveal 824 transfer events and 5,330 sale events (accounting for a total of \$8,857,070.41) and 370 address pairs related to NFT wash trading behaviors, causing a minimum loss of \$3,965,247.13. Furthermore, we provide insights from six aspects, i.e., marketplace design, profitability, NFT project design, payment token, user behavior, and NFT ecosystem.
Index Terms:
NFT, OpenSea, Blockchain, Wash Trading, Empirical Study, Transaction Analysis
I Introduction
@@
NFTs (Non-Fungible Tokens) [1] are blockchain-enabled digital assets, which users can buy and sell without third-party participation. Due to the rising enthusiasm for the concept, the NFT trading volume increased to an astonishing \$25.1 billion in 2021 [2]. Many NFT marketplaces have been created to facilitate NFT trading. Take OpenSea, the NFT marketplace with over one million registered users [3], as an example. It facilitated transactions (TXNs) for around \$5 billion alone in January 2022 [4]. However, the market value of NFTs may not justify such a thriving market. According to Binance, almost 45% of all NFT trading volume may be fraudulent due to wash trading events [5], where users manipulate the market by buying and selling the same financial product [6]. A notorious case  [7, 8, 9, 10, 11] involves the wash trading on Cryptopunk #9998 (an NFT) [^1] The reference of all the NFTs, TXN hashes, and Ethereum accounts mentioned in each section can be found at REFERENCE.csv, where the buyer initially obtained loans from multiple sources to purchase the NFT, then immediately sold the NFT to the origin holder for the same price, and finally repaid the loans. Another example is that Meebits #13824 was traded twice between 0x35D0CA and 0xA99A76 for 14,700 WETH [12](a cryptocurrency) and 15,000 WETH, around 40 times the previous trading price.
@@
The existence of NFT wash trading has been proven in the previous works [13, 14, 15, 16, 17], with several features abstracted, e.g., Von et al. [13] discovered that 2.04% of NFTs’ sale events trigger suspicions of market abuse, while Das et al. [14] adopted the strongly/weakly connected component of the graph constructed by NFTs’ events to detect NFT wash trading. However, they mainly abstracted NFT wash trading as graph patterns and only focused on its financial losses, while ignoring TXNs themselves (e.g., the trading price) and other aspects of the NFT ecosystem. Investigating NFT wash trading remains challenging due to the lack of a broader discussion on its definition, a researcher-friendly dataset, and more valuable insights. To fill the gap, we present the most comprehensive study on NFT wash trading, with findings from six aspects.
@@
In this paper, we first collect NFTs’ sale/transfer events through API access and related block/ERC-20 TXNs via open-source datasets. Then, we uncover and address three previously unnoticed issues about the validity of NFTs’ data via the OpenSea API (a widely used API for collecting NFT’s trading information [18]), i.e., inconsistency in NFTs’ event sequences, timestamp errors in records, and mislabeling of sale/transfer events. Furthermore, we define three types of NFT wash trading, i.e., Round-trip Trading, Unprofitable Trading, and Hidden Trading. Then, we design heuristic algorithms to detect each type and adopt FP-Growth [19] to identify wash trading address pairs/groups. We flag 824 transfer events, 5,330 sale events, 370 address pairs, and 29 address groups related to wash trading, accounting for tokens worth around \$8,857,070.41. Based on the experimental results, we offer insights from six aspects, i.e., marketplace design, profitability, NFT project design, payment token, user behavior, and the NFT ecosystem. The main contributions of this paper are summarized as follows:
@@
  • We contribute the most extensive dataset of 2,701,883 NFTs’ event sequences from 285 most popular collections. We are the first study to raise and address problems of NFT data collected by OpenSea API. We have released the open-source raw and processed datasets to help researchers uncover further studies222 https://github.com/NFTWashTrading/The_Dark_Side_of_NFTs.
@@ II-A Ethereum
Based on blockchain technology, Ethereum is a distributed ledger platform with programmable features. With the design of smart contracts and the account-based model, Ethereum allows people to initiate TXNs and develop applications for users to interact with on blockchains [20]. Smart contracts. Smart contracts [21] are tamper-proof programs stored on the blockchain that run when predetermined conditions are met. They facilitate TXNs in the decentralized system. Account. Ethereum has two account types, i.e., externally-owned account (EOA) and contract account [22]. EOAs are controlled by anyone with private keys, while the contract account is associated with the smart contract code [23]. Ether and ERC-20 token. Ether (ETH) is the native token circulating in Ethereum, used as a payment system for verifying TXNs [20]. ERC-20 token is a standard for creating alternative tokens developed by smart contracts [24]. Block (External) TXN, internal TXN, and ERC-20 token TXN. A block TXN [25] constructs the body of the block in Ethereum and refers to the TXN where the sender sends ETH or other tokens to the receiver with some additional information (e.g., smart contract function calls). An internal TXN refers to the TXN that occurs during the execution of a smart contract [26]. When a smart contract is called, it may execute multiple functions, and each function may trigger calls to other contracts. The invocation and interaction between these contracts are achieved through internal TXNs. In this paper, we refer to all TXNs involving the transfer of ERC-20 tokens between addresses as ERC-20 token TXNs.
@@
Figure 1: Two main ways to purchase NFTs on OpenSea
II-B NFT Trading
@@
An NFT (Non-Fungible Token) is a unique on-chain identifier that typically refers to digital files. The tokenId and contract address represent a unique NFT on the blockchain. The first transfer event of an NFT is also called the minted event, showing the publishing of an NFT on the blockchain. NFTs are often purchased on various trading platforms, such as OpenSea. Regarding OpenSea, as shown in Figure 1, there are two main ways to purchase an NFT. 1) instant sale (Black), where an NFT is listed at a fixed price, and a buyer can purchase it immediately, with a portion of the trading value being transferred to OpenSea’s official wallet as service fees. 2) auction(Green), where buyers place bids on an NFT, and sellers can accept the offer. In this case, the sellers pay the service fees (part of the auction price) to OpenSea. The payment token for the auction is WETH [12].
III WASH TRADING/TRADER TYPES
In this section, we discuss the possible forms of NFT wash trading and wash traders, which have not been mentioned in previous works [13, 14, 27], with detailed examples.
III-A Type 1: Round-trip Trading
@@
Explanation. Round-trip Trading is the unethical practice of repeatedly buying and selling the same securities to manipulate the finance market [28]. In the NFT market, Round-trip Trading happens when someone purchases an NFT and promptly resells it, either directly or through multiple addresses, to the original NFT holder. Example. OG:Crystal #4015 was traded 24 times at almost the same price between 0xCF6FF6 and 0xC17D7c within 9 hours, accounting for 126.2 ETH.
III-B Type 2: Unprofitable Trading
@@
Explanation. We define that Unprofitable Trading refers to an NFT TXN where the buyer is either funded beforehand by the seller or receives the seller’s return amount, indicating that the TXN is not intended for profit. Normally, the transfer of NFT (seller $\rightarrow$ buyer) and the transfer of TXN amount (buyer $\rightarrow$ seller) occur simultaneously within a block TXN, which is guaranteed by smart contracts. However, Unprofitable Trading occurs with an additional value transfer, e.g., the seller transfers a certain amount of ETH that is similar to the NFT trading price to the buyer shortly before or after the TXN. Another conduction is using a certain amount of ERC-20 tokens with market value instead of ETH. An early value transfer can be interpreted as a form of funding, while the restitution of the amount illustrates that the NFT TXN does not profit from a third party. Example. For Omnimorph #3980, 0x6149ca (seller) transferred 0.23 ETH (1.045% of the trading price) to 0x2beba3 (buyer) three minutes before the NFT sale event. Twenty minutes later, 0x2beba3 (new seller) returned 0.2098 ETH (0.95% of the trading price) to 0x0f767ef (new buyer) four minutes after the next sale event. The two NFT sellers did not profit from either TXN, which was abnormal for a usual sale event. They conspired to complete the wash trading process.
III-C Type 3: Hidden Trading
@@
Explanation. Hidden Trading is third type of wash trading we identify, characterized by collusive actions between sellers and designated buyers. Hidden Trading implies the existence of a series of continuous private tradings where the NFT sellers designate the NFT buyers. On OpenSea, NFT holders can reserve items for specific buyers [29], i.e., only the addresses approved by the holders have the right to purchase the NFT. Private Tradings disregard the impact of market liquidity and user sentiment because the sellers and buyers know each other before the TXNs occur. Example. Wash traders raised the NFT price from 3.750 ETH, 3.780 ETH to 4 ETH, 7.4 ETH for VeeFriends #7582, all through continuous private tradings.
III-D Wash trading pairs/groups
@@
IV DATA COLLECTION AND PREPROCESSING
In this section, we detail the methodology used to construct our dataset, which is derived from API access and open-source datasets. Subsequently, we reveal and address three problems associated with the data collected via OpenSea API.
IV-A Data collection
@@
TABLE I: An example of event sequence: Alpha Shark #9’s event sequence.
IV-A1 Event sequence
@@
The OpenSea API allows users to fetch metadata and core elements of NFTs [18], e.g., sale/transfer events. The transfer events for each NFT reveal its circulation flow, while the sale events additionally provide specific trading information. We retrieve 8,717,031 transfer events and 3,830,141 sale events of 2,701,883 NFTs from Ethereum’s 285 most popular collections. We start counting each NFT’s event sequence since it was minted (00th record), and the $i$th record’s fields will be marked as timestamp_i, etc. Table I displays an example of the event sequence, i.e., since Alpha Shark (collection) #9 (tokenId) was minted, it was transferred to 0x1c2fd0 (to_0), then to 0x9164e3 (to_1), then bought by 0x99264d (to_2) for 2.9 (numToken_2) ETH (payToken_2) at 2022-06-15T16:56:53 (timestamp_2), while each ETH cost \$1215.68 (usdToken_2). Also, isPrivate indicates whether the NFT is reserved for a specific buyer.
IV-A2 Block TXN and ERC-20 token TXN
@@
XBlock [30] is a data platform for the blockchain community. This data source [25] provides researchers with information about block TXNs and ERC-20 token TXNs. To enable the identification of ETH and ERC-20 token transfers, we filter out 184,008,844 block TXNs and 48,513,194 ERC-20 token TXNs in the field of the same from or to as that of each NFT sale/transfer event.
IV-A3 Historical market price data
@@
CoinGecko API enables users to obtain crypto prices, historical market data, etc [31]. We adopt it to collect the historical market price of ERC-20 tokens. Specifically, we identify 233,618 ERC-20 token smart contracts involved in related ERC-20 token TXNs. Since not all ERC-20 tokens have a market price reference value, we finally collect 2,373,787 historical price records of 2,982 ERC-20 tokens.
IV-B PREPROCESSING
We primarily address three issues related to NFT event sequences collected via the OpenSea API and propose corresponding processing rules to enhance data validity.
IV-B1 Address inconsistency
[…]
timestamp tokenId from to type
2021-08-15T16:03:52 305 0x000000 0x01c9e1 transfer
2021-08-15T16:05:05 305 0x000000 0x01c9e1 transfer
TABLE II: An example of the repeated minted event: 0n1-force #305 with two minted events in the event sequence.
IV-B2 Timestamp error
Issue. The timestamps in event sequences are not always correct, e.g., the timestamp of TXN 0x2c8097 is 1,648,160,557. However, the actual timestamp on the blockchain is 1,648,160,556. We assume the timestamp error might be the technical failures of the OpenSea itself, as we find some online complaints about the slow and failed metadata updates of OpenSea [32, 33]. It suggests that OpenSea’s data synchronization is in error to some extent. Preprocessing. We discard all events with incorrect timestamps as the timestamp is a critical factor in segmenting time windows in our algorithm (See V-A2 for details). It turns out that records with incorrect timestamps account for less than 0.001%, having little impact on our later work.
IV-B3 Fake events
Issue. It is not easy to completely and correctly label NFT’s transfer/sale events. We explore the following reasons. 1) A sale event displayed on OpenSea’s UI or retrieved via API is a TXN processed by one of the marketplace smart contracts (e.g., OpenSea: Wyvern Exchange v1) shown on Etherscan [34]. However, transfer events could essentially be sale events as new NFT marketplaces emerge. Meanwhile, their TXNs are not logged synchronously into OpenSea. For example, TXN 0x361ae0 is marked as a transfer event but shows an NFT trading behavior where Loot #2157 was sold for 0.997 ETH on  [35]Element (a marketplace). It may lead to some frequent trading being treated as multiple free transfers and therefore being overlooked as suspicious. 2) Marketplace contracts might appear as to or from in records, misleading the trace of an NFT. For instance, say user A invokes a marketplace contract named GemSwap2 to buy an NFT from user B. In OpenSea, the first record is a sale event where A first sales the NFT to GemSwap2, then followed by a transfer event in which GemSwap2 transfers it to B. GemSwap2 is redundant as an intermediate address and may cause confusion by appearing as a frequent NFT trader. Preprocessing. We ignore the potential sale events since we also include transfer events to identify wash trading behaviors. Then, we merge all the redundant intermediate smart contracts related to marketplace contracts.
IV-C Dataset overview

### v2

@@ Abstract.
NFTs (Non-Fungible Tokens) have seen significant growth since they first captured public attention in 2021. However, the NFT market is plagued by fake transactions and economic bubbles, e.g., NFT wash trading. Wash trading typically refers to a transaction involving the same person or two colluding individuals, and has become a major threat to the NFT ecosystem. Previous studies only detect NFT wash trading from the financial aspect, while the real-world wash trading cases are much more complicated (e.g., not aiming at inflating the market value). There is still a lack of multi-dimension analysis to better understand NFT wash trading. Therefore, we present the most comprehensive study of NFT wash trading, analyzing 8,717,031 transfer events and 3,830,141 sale events from 2,701,883 NFTs. We identify three types of NFT wash trading and propose identification algorithms. Our experimental results reveal 824 transfer events and 5,330 sale events (accounting for a total of \$8,857,070.41) and 370 address pairs related to NFT wash trading behaviors, causing a minimum loss of \$3,965,247.13. Furthermore, we provide insights from six aspects, i.e., marketplace design, profitability, NFT project design, payment token, user behavior, and NFT ecosystem.
Blockchain, Ethereum, Non-Fungible Tokens, Wash Trading
journalyear: 2024copyright: acmlicensedconference: 15th Asia-Pacific Symposium on Internetware; July 24–26, 2024; Macau, Macaobooktitle: 15th Asia-Pacific Symposium on Internetware (Internetware 2024), July 24–26, 2024, Macau, Macaodoi: 10.1145/3671016.3674808isbn: 979-8-4007-0705-6/24/07ccs: Security and privacy Human and societal aspects of security and privacy
1. Introduction
@@
NFTs (Non-Fungible Tokens) (Wikipedia, 2023) are blockchain-enabled digital assets, which users can buy and sell without third-party participation. Due to the rising enthusiasm for the concept, the NFT trading volume increased to an astonishing \$25.1 billion in 2021 (Hayward, 2022). Many NFT marketplaces have been created to facilitate NFT trading. Take OpenSea, the NFT marketplace with over one million registered users (Hayes, 2022), as an example. It facilitated transactions (TXNs) for around \$5 billion alone in January 2022 (Hypebeast, 2023). However, the market value of NFTs may not justify such a thriving market. According to Binance, almost 45% of all NFT trading volume may be fraudulent due to wash trading events (Coinfomania, 2022), where users manipulate the market by buying and selling the same financial product (Sergeenkov, 2022). A notorious case  (Tahmasbi and Fuchsberger, 2022; Staff, 2021; pastel, 2022; Baker, 2021; ANDREW ROSSOW, 2022) involves the wash trading on Cryptopunk #9998 (an NFT) [^1] The reference of all the NFTs, TXN hashes, and Ethereum accounts mentioned in each section can be found at REFERENCE.csv, where the buyer initially obtained loans from multiple sources to purchase the NFT, then immediately sold the NFT to the origin holder for the same price, and finally repaid the loans. Another example is that Meebits #13824 was traded twice between 0x35D0CA and 0xA99A76 for 14,700 WETH (Wra, 2023)(a cryptocurrency) and 15,000 WETH, around 40 times the previous trading price.
@@
The existence of NFT wash trading has been proven in the previous works (von Wachter et al., 2022; Das et al., 2021; Serneels, 2023; Tariq and Sifat, 2022; Wen et al., 2023), with several features abstracted, e.g., von Wachter et al. (von Wachter et al., 2022) discovered that 2.04% of NFTs’ sale events trigger suspicions of market abuse, while Das et al. (Das et al., 2021) adopted the strongly/weakly connected component of the graph constructed by NFTs’ events to detect NFT wash trading. However, they mainly abstracted NFT wash trading as graph patterns and only focused on its financial losses, while ignoring TXNs themselves (e.g., the trading price) and other aspects of the NFT ecosystem. Investigating NFT wash trading remains challenging due to the lack of a broader discussion on its definition, a researcher-friendly dataset, and more valuable insights. To fill the gap, we present the most comprehensive study on NFT wash trading, with findings from six aspects.
@@
In this paper, we first collect NFTs’ sale/transfer events through API access and related block/ERC-20 TXNs via open-source datasets. Then, we define three types of NFT wash trading, i.e., Round-trip Trading, Unprofitable Trading, and Hidden Trading. Next, we design heuristic algorithms to detect each type and adopt FP-Growth (Han et al., 2004) to identify wash trading address pairs/groups. We flag 824 transfer events, 5,330 sale events, 370 address pairs, and 29 address groups related to wash trading, accounting for tokens worth around \$8,857,070.41. Based on the experimental results, we offer insights from six aspects, i.e., marketplace design, profitability, NFT project design, payment token, user behavior, and the NFT ecosystem. The main contributions of this paper are summarized as follows:
@@
  • We contribute an extensive dataset of 2,701,883 NFTs’ event sequences from 285 most popular collections. We have released the open-source processed datasets to help researchers uncover further studies222 https://github.com/NFTWashTrading/The_Dark_Side_of_NFTs.
@@ 2.1. Ethereum
Based on blockchain technology, Ethereum is a distributed ledger platform with programmable features. With the design of smart contracts and the account-based model, Ethereum allows people to initiate TXNs and develop applications for users to interact with on blockchains (wik, 2023a). Smart contracts (wik, 2023b) are tamper-proof programs stored on the blockchain that run when predetermined conditions are met. They facilitate TXNs in the decentralized system. Ethereum has two account types, i.e., externally-owned account (EOA) and contract account (contributors, 2023). EOAs are controlled by anyone with private keys, while the contract account is associated with the smart contract code (Vujičić et al., 2018). Ether (ETH) is the native token circulating in Ethereum, used as a payment system for verifying TXNs (wik, 2023a). ERC-20 token is a standard for creating alternative tokens developed by smart contracts (OpenZeppelin, 2023). A block TXN (Zheng et al., 2020) constructs the body of the block in Ethereum and refers to the TXN where the sender sends ETH or other tokens to the receiver with some additional information (e.g., smart contract function calls). An internal TXN refers to the TXN that occurs during the execution of a smart contract (Chan and Olmsted, 2017). When a smart contract is called, it may execute multiple functions, and each function may trigger calls to other contracts. The invocation and interaction between these contracts are achieved through internal TXNs. In this paper, we refer to all TXNs involving the transfer of ERC-20 tokens between addresses as ERC-20 token TXNs.
@@
Figure 1. Two main ways to purchase NFTs on OpenSea: instant sale (Black), auction (Green).
2.2. NFT Trading
@@
An NFT (Non-Fungible Token) is a unique on-chain identifier that typically refers to digital files. The tokenId and contract address represent a unique NFT on the blockchain. The first transfer event of an NFT is also called the minted event, showing the publishing of an NFT on the blockchain. NFTs are often purchased on various trading platforms, such as OpenSea. Regarding OpenSea, as shown in Figure 1, there are two main ways to purchase an NFT. 1) instant sale, where an NFT is listed at a fixed price, and a buyer can purchase it immediately, with a portion of the trading value being transferred to OpenSea’s official wallet as service fees. 2) auction, where buyers place bids on an NFT, and sellers can accept the offer. In this case, the sellers pay the service fees (part of the auction price) to OpenSea. The payment token for the auction is WETH (Wra, 2023).
3. WASH TRADING/TRADER TYPES
3.1. Type 1: Round-trip Trading
@@
Explanation. Round-trip Trading is the unethical practice of repeatedly buying and selling the same securities to manipulate the finance market (SCOTT, 2023). In the NFT market, Round-trip Trading happens when someone purchases an NFT and promptly resells it, either directly or through multiple addresses, to the original NFT holder. Example. OG:Crystal #4015 was traded 24 times at almost the same price between 0xCF6FF6 and 0xC17D7c within 9 hours, accounting for 126.2 ETH.
3.2. Type 2: Unprofitable Trading
@@
Explanation. We define that Unprofitable Trading refers to an NFT TXN where the buyer is either funded beforehand by the seller or receives the seller’s return amount, indicating that the TXN is not intended for profit. Normally, the transfer of NFT (seller $\rightarrow$ buyer) and the transfer of TXN amount (buyer $\rightarrow$ seller) occur simultaneously within a block TXN, which is guaranteed by smart contracts. However, Unprofitable Trading occurs with an additional value transfer, e.g., the seller transfers a certain amount of ETH that is similar to the NFT trading price to the buyer shortly before or after the TXN. Another conduction is using a certain amount of ERC-20 tokens with market value instead of ETH. An early value transfer can be interpreted as a form of funding, while the restitution of the amount illustrates that the NFT TXN does not profit from a third party. Example. For Omnimorph #3980, 0x6149ca (seller) transferred 0.23 ETH (104.5% of the trading price) to 0x2beba3 (buyer) three minutes before the NFT sale event. Twenty minutes later, 0x2beba3 (new seller) returned 0.2098 ETH (95% of the trading price) to 0x0f767ef (new buyer) four minutes after the next sale event. The two NFT sellers did not profit from either TXN, which was abnormal for a usual sale event. They conspired to complete the wash trading process.
3.3. Type 3: Hidden Trading
@@
Explanation. Hidden Trading is third type of wash trading we identify, characterized by collusive actions between sellers and designated buyers. Hidden Trading implies the existence of a series of continuous private tradings where the NFT sellers designate the NFT buyers. On OpenSea, NFT holders can reserve items for specific buyers (Goodman, 2022), i.e., only the addresses approved by the holders have the right to purchase the NFT. Private Tradings disregard the impact of market liquidity and user sentiment because the sellers and buyers know each other before the TXNs occur. Example. Wash traders raised the NFT price from 3.750 ETH, 3.780 ETH to 4 ETH, 7.4 ETH for VeeFriends #7582, all through continuous private tradings.
3.4. Wash trading pairs/groups
@@
4. DATA COLLECTION
@@
Table 1. An example of event sequence: Alpha Shark #9’s event sequence.
4.1. Event sequence
@@
The OpenSea API allows users to fetch metadata and core elements of NFTs (OpenSea, 2023), e.g., sale/transfer events. The transfer events for each NFT reveal its circulation flow, while the sale events additionally provide specific trading information. We retrieve 8,717,031 transfer events and 3,830,141 sale events of 2,701,883 NFTs from Ethereum’s 285 most popular collections. We start counting each NFT’s event sequence since it was minted (00th record), and the $i$th record’s fields will be marked as timestamp_i, etc. Table 1 displays an example of the event sequence, i.e., since Alpha Shark (collection) #9 (tokenId) was minted, it was transferred to 0x1c2fd0 (to_0), then to 0x9164e3 (to_1), then bought by 0x99264d (to_2) for 2.9 (numToken_2) ETH (payToken_2) at 2022-06-15T16:56:53 (timestamp_2), while each ETH cost \$1215.68 (usdToken_2). Also, isPrivate indicates whether the NFT is reserved for a specific buyer.
4.2. Block TXN and ERC-20 token TXN
@@
XBlock (xbl, 2023) is a data platform for the blockchain community. This data source (Zheng et al., 2020) provides researchers with information about block TXNs and ERC-20 token TXNs. To enable the identification of ETH and ERC-20 token transfers, we filter out 184,008,844 block TXNs and 48,513,194 ERC-20 token TXNs in the field of the same from or to as that of each NFT sale/transfer event.
4.3. Historical market price data
@@
CoinGecko API enables users to obtain crypto prices, historical market data, etc (CoinGecko, 2023). We adopt it to collect the historical market price of ERC-20 tokens. Specifically, we identify 233,618 ERC-20 token smart contracts involved in related ERC-20 token TXNs. Since not all ERC-20 tokens have a market price reference value, we finally collect 2,373,787 historical price records of 2,982 ERC-20 tokens.
4.4. Dataset overview